i need help with a math problem i need it to be solved
the problem is (40-2x)(17-2x) i know that the answer is 222 but i don't know how that is possible

Answer by stanbon(48568)

(40-2x)(17-2x)
Use FOIL:
F: 40*17 = 680
O: -2x*40 = -80x
I: -2x*17 = -34x
L: -2x*-2x = 4x^2
------------------
Combine the terms to get:
4x^2-114x+680
==============
Cheers,
Stan H.